# Intrinsic back-switching phenomenon in SOT-MRAM devices

### v2

@@ IV Discussion and Simulations
To go further in the analysis of these results, we compared them with the results obtained from simulations. It is known that for pillar diameters such as those we have measured, the reversal of magnetization by SOT occurs through a two-step nucleation-propagation process [34, 35]. Micromagnetic simulations are therefore the most accurate way of faithfully reproducing magnetization behavior. However, such simulations to calculate WER curves are time-consuming. Therefore, we aimed to reproduce these measurements using macrospin simulations. In this coherent reversal approximation, the magnetization dynamics can be reproduced by the Landau-Lifshitz-Gilbert equation, to which the two terms describing the FL and DL torques of the SOTs [20] are added:
@@
is obtained by Vibrating Sample Magnetometer measurements, neglecting the effect of any dead layer. , , and are imposed. is initially taken to be 0.033 [36, 37] and and H are first roughly estimated by fitting the SOT current vs H forward switching curve using the equation established in [19]:
@@
The comparison of the results of the numerical simulations and those of the experiments is shown in Fig. 3 (d), (e) and (f). The simulations do not reproduce all the characteristics of the experimental curves, for example, the variation of the switching current as a function of H field, which is very abrupt in the simulations compared with the behavior obtained experimentally. Nevertheless, the general agreement is very good, and the different regions of no-switching, deterministic switching with minimal WER, and loss of determinism in the switching, are reproduced very well. Furthermore, the parameter values used to obtain this agreement are realistic and very close to those reported in the literature for W/CoFeB/MgO multilayers (see for example [2] and references herein). This is particularly noteworthy given that, as explained above, the dynamics of magnetization reversal by SOT can only be described in detail using a micromagnetic model.
This agreement between experimental and simulated results could point to the predominant role of nucleation in determining threshold currents for the boundaries between no-switching/deterministic switching and deterministic switching/back-switching for sub- nm SOT-MRAM devices [25].
@@
In the presence of an applied in-plane field, H, the two equilibrium positions m are shifted towards the x-direction such that their new coordinates can be given by m and m where , as explained in [22]. In our simulations for I and H, the initial state of magnetization is m-. During the application of the current pulse, the magnetization reaches a state of dynamic equilibrium under the predominant action of SOT, the applied magnetic field and the anisotropy. At the end of the pulse, the magnetization relaxes under the effect of the applied magnetic field, the anisotropy and thermal fluctuations.
@@
When the current is increased, in the region of deterministic magnetization reversal, the current-induced torques are strong enough to drive the magnetization to a dynamic equilibrium close to m+ with small negative x and positive y components. At the end of the current pulse, the magnetization then relaxes to the m+ state, resulting in a deterministic switching, as shown in Fig. 4 (a) for IA and H Oe (green curve). By further increasing the current, the dynamic equilibrium begins to shift towards a fully in-plane position. This position is energetically unstable, and the resulting final state after the relaxation is highly sensitive to any energy or time fluctuation, as reported in [22]. As shown in Fig. 4 (a) (blue curve) for IA and H Oe, the magnetization can then return to its original m- state, resulting in a BSW phenomenon.
@@
With the same arguments, we can understand the impact of H by comparing the switching orbits for IA and H Oe in the presence of a magnetic field H Oe and Oe as shown in Fig. 4 (c). By expanding equation 1, we obtain the component along the y axis of the SOT effective field: H [27]. With the values given in Table 1, H is of the same sign as H. H is therefore negative for a positive injected current. Starting from H (Fig. 4 (a) blue curves), the application of H (H Oe) will reinforce the negative component of the effective field along y and further bring the dynamic equilibrium in-plane, reinforcing the non-deterministic character of the reversal, as shown by the yellow curve in Fig. 4 (c). H (H Oe), on the other hand, will oppose H and move the dynamic equilibrium away from this energetically unstable position. Consequently, the application of H Oe leads to a deterministic reversal, as shown by the brown curve in Fig. 4 (c). It should be noted that these results are in line with previous studies showing deterministic switching for and non-deterministic switching for [13, 14, 25]. In the second case, the two y-components of the SOT effective fields will add up to bring the dynamic equilibrium to an in-plane position.
@@
Since the final state depends strongly on the relaxation process, it should also depend on the magnetic damping parameter as reported in [19, 20, 24]. Fig. 4 (d) shows the switching orbits for IA, H Oe and different . The damping coefficient , which we used in our simulations, results in slow energy dissipation and hence randomness in the final state (blue curve). This randomness is even more prevalent when is lower, as shown by the light blue curve for . As increases, energy is dissipated more rapidly and randomness is reduced, eventually disappearing, as shown by the red curve for and as reported previously [24]. Unlike STT switching current, which scales with , SOT switching current is little affected by a large magnetic damping parameter [19, 30]. Increasing the magnetic damping parameter can therefore significantly reduce the BSW without severely affecting the SOT-MRAM write current. Finally, since thermal fluctuations introduce randomness in the current-induced torques during the relaxation process, higher temperatures show an increase in WER. At low temperature, the thermal fluctuations are minimal and the final state depends only on the dynamic equilibrium and .
@@ V Reducing the backswitching by shaping the current pulse
There is another way of reducing BSW and improving WER without having to change the material system. It consists of increasing the length of the fall time of the write current pulse, as predicted numerically [23]. Fig. 5 (a) shows WER curves obtained experimentally in the presence of an in-plane field, H Oe and injecting 10 ns-long SOT current pulses. A widening of the forward switching region is observed as the fall time increases from 2 ns to 4 ns, with the BSW phenomenon disappearing for the current range used in the case of a fall time of 4 ns. This can be understood using the same arguments as above, by looking at magnetization relaxation at the end of the pulse. When the fall time is non-zero, current-induced torques are present during relaxation. The longer the fall time, the longer their amplitude remains at high values. The impact of thermal fluctuations is then reduced, but more importantly, the magnetization is gradually brought out of an energetically unstable position. In this way, the dynamic equilibrium is gradually modified. Such behavior is in line with the explanation given in [23]. However, much longer fall times (4 ns) were required in our experiments compared with those obtained in the simulations of [23], typically of the order of 0.2 ns. It may be attributed to thermal fluctuations and Joule heat effects, which are not taken into account in the simulations. This difference needs to be investigated. In any case, the results of our experiments show a simple solution to the BSW phenomenon that can be easily implemented without the need for further material developments. We tested this solution on a complete SOT-MRAM single cell (Fig. 5 (b)). We improved the WER by almost 3 orders of magnitude using a longer fall time of 2 ns, and obtained no error for write attempts.
@@ Acknowledgements.
We thank the staff of Antaios for their support throughout this project. We thank Chando Park, Minrui Yu, Luc Thomas and Mahendra Pakala from Applied Materials for fruitful discussions. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 955671, and was supported by the Région Auvergne-Rhône-Alpes Pack Ambition Recherche Program (grant 19-009938-01-MAPS) and the French RENATECH network implemented at the Upstream Technological Platform in Grenoble PTA (ANR-22-PEEL-0015).References
- Prenat et al. [2016] G. Prenat, K. Jabeur, P. Vanhauwaert, G. D. Pendina, F. Oboril, R. Bishnoi, M. Ebrahimi, N. Lamard, O. Boulle, K. Garello, J. Langer, B. Ocker, M.-C. Cyrille, P. Gambardella, M. Tahoori, and G. Gaudin, Ultra-Fast and High-Reliability SOT-MRAM: From Cache Replacement to Normally-Off Computing, IEEE Transactions on Multi-Scale Computing Systems 2, 49 (2016).
- Shao et al. [2021] Q. Shao, P. Li, L. Liu, H. Yang, S. Fukami, A. Razavi, H. Wu, K. Wang, F. Freimuth, Y. Mokrousov, M. D. Stiles, S. Emori, A. Hoffmann, J. Akerman, K. Roy, J.-P. Wang, S.-H. Yang, K. Garello, and W. Zhang, Roadmap of Spin–Orbit Torques, IEEE Transactions on Magnetics 57, 1 (2021).
- [3] Spin-transfer Torque DDR Products, Everspin, https://www.everspin.com/spin-transfer-torque-ddr-products, accessed on 2023-04-12.
- Miron et al. [2011] I. M. Miron, K. Garello, G. Gaudin, P.-J. Zermatten, M. V. Costache, S. Auffret, S. Bandiera, B. Rodmacq, A. Schuhl, and P. Gambardella, Perpendicular switching of a single ferromagnetic layer induced by in-plane current injection, Nature 476, 189 (2011).
- Cubukcu et al. [2014] M. Cubukcu, O. Boulle, M. Drouard, K. Garello, C. Onur Avci, I. Mihai Miron, J. Langer, B. Ocker, P. Gambardella, and G. Gaudin, Spin-orbit torque magnetization switching of a three-terminal perpendicular magnetic tunnel junction, Applied Physics Letters 104, 042406 (2014).
- Liu et al. [2012] L. Liu, C.-F. Pai, Y. Li, H. W. Tseng, D. C. Ralph, and R. A. Buhrman, Spin-Torque Switching with the Giant Spin Hall Effect of Tantalum, Science 336, 555 (2012).
- You et al. [2015] L. You, O. Lee, D. Bhowmik, D. Labanowski, J. Hong, J. Bokor, and S. Salahuddin, Switching of perpendicularly polarized nanomagnets with spin orbit torque without an external magnetic field by engineering a tilted anisotropy, Proceedings of the National Academy of Sciences 112, 10310 (2015).
- Cai et al. [2021] W. Cai, K. Shi, Y. Zhuo, D. Zhu, Y. Huang, J. Yin, K. Cao, Z. Wang, Z. Guo, Z. Wang, G. Wang, and W. Zhao, Sub-ns Field-Free Switching in Perpendicular Magnetic Tunnel Junctions by the Interplay of Spin Transfer and Orbit Torques, IEEE Electron Device Letters 42, 704 (2021).
- van den Brink et al. [2016] A. van den Brink, G. Vermijs, A. Solignac, J. Koo, J. T. Kohlhepp, H. J. M. Swagten, and B. Koopmans, Field-free magnetization reversal by spin-Hall effect and exchange bias, Nature Communications 7, 10854 (2016).
- Garello et al. [2019] K. Garello, F. Yasin, H. Hody, S. Couet, L. Souriau, S. H. Sharifi, J. Swerts, R. Carpenter, S. Rao, W. Kim, J. Wu, K. Sethu, M. Pak, N. Jossart, D. Crotti, A. Furnémont, and G. S. Kar, Manufacturable 300mm platform solution for Field-Free Switching SOT-MRAM, in 2019 Symposium on VLSI Technology (2019) pp. T194–T195.
- Kao et al. [2022] I.-H. Kao, R. Muzzio, H. Zhang, M. Zhu, J. Gobbo, S. Yuan, D. Weber, R. Rao, J. Li, J. H. Edgar, J. E. Goldberger, J. Yan, D. G. Mandrus, J. Hwang, R. Cheng, J. Katoch, and S. Singh, Deterministic switching of a perpendicularly polarized magnet using unconventional spin–orbit torques in WTe2, Nature Materials 21, 1029 (2022).
- Krizakova et al. [2022a] V. Krizakova, M. Perumkunnil, S. Couet, P. Gambardella, and K. Garello, Spin-orbit torque switching of magnetic tunnel junctions for memory applications, Journal of Magnetism and Magnetic Materials 562, 169692 (2022a).
- Lee et al. [2018] J. M. Lee, J. H. Kwon, R. Ramaswamy, J. Yoon, J. Son, X. Qiu, R. Mishra, S. Srivastava, K. Cai, and H. Yang, Oscillatory spin-orbit torque switching induced by field-like torques, Communications Physics 1, 2 (2018).
- Yoon et al. [2017] J. Yoon, S.-W. Lee, J. H. Kwon, J. M. Lee, J. Son, X. Qiu, K.-J. Lee, and H. Yang, Anomalous spin-orbit torque switching due to field-like torque–assisted domain wall reflection, Science Advances 3, e1603099 (2017).
- Decker et al. [2017] M. Decker, M. Wörnle, A. Meisinger, M. Vogel, H. Körner, G. Shi, C. Song, M. Kronseder, and C. Back, Time Resolved Measurements of the Switching Trajectory of Pt/Co Elements Induced by Spin-Orbit Torques, Physical Review Letters 118, 257201 (2017).
- Min et al. [2009] T. Min, J. Z. Sun, R. Beach, D. Tang, and P. Wang, Back-hopping after spin torque transfer induced magnetization switching in magnetic tunneling junction cells, Journal of Applied Physics 105, 07D126 (2009).
- Kim et al. [2016] W. Kim, S. Couet, J. Swerts, T. Lin, Y. Tomczak, L. Souriau, D. Tsvetanova, K. Sankaran, G. L. Donadio, D. Crotti, S. Van Beek, S. Rao, L. Goux, G. S. Kar, and A. Furnemont, Experimental Observation of Back-Hopping With Reference Layer Flipping by High-Voltage Pulse in Perpendicular Magnetic Tunnel Junctions, IEEE Transactions on Magnetics 52, 1 (2016).
- Devolder et al. [2020] T. Devolder, O. Bultynck, P. Bouquin, V. D. Nguyen, S. Rao, D. Wan, B. Sorée, I. P. Radu, G. S. Kar, and S. Couet, Back hopping in spin transfer torque switching of perpendicularly magnetized tunnel junctions, Physical Review B 102, 184406 (2020).
- Lee et al. [2013] K.-S. Lee, S.-W. Lee, B.-C. Min, and K.-J. Lee, Threshold current for switching of a perpendicular magnetic layer induced by spin Hall effect, Applied Physics Letters 102, 112410 (2013).
- Park et al. [2014] J. Park, G. E. Rowlands, O. J. Lee, D. C. Ralph, and R. A. Buhrman, Macrospin modeling of sub-ns pulse switching of perpendicularly magnetized free layer via spin-orbit torques for cryogenic memory applications, Applied Physics Letters 105, 102404 (2014).
- Legrand et al. [2015] W. Legrand, R. Ramaswamy, R. Mishra, and H. Yang, Coherent Subnanosecond Switching of Perpendicular Magnetization by the Fieldlike Spin-Orbit Torque without an External Magnetic Field, Physical Review Applied 3, 064012 (2015).
- Taniguchi [2019] T. Taniguchi, Theoretical condition for switching the magnetization in a perpendicularly magnetized ferromagnet via the spin Hall effect, Physical Review B 100, 174419 (2019).
- Zhu and Zhao [2020] D. Zhu and W. Zhao, Threshold Current Density for Perpendicular Magnetization Switching Through Spin-Orbit Torque, Physical Review Applied 13, 044078 (2020).
- Taniguchi et al. [2020] T. Taniguchi, Y. Shiokawa, and T. Sasaki, Reduction of back switching by large damping ferromagnetic material, Applied Physics Express 13, 123002 (2020).
- Krizakova et al. [2022b] V. Krizakova, M. Hoffmann, V. Kateel, S. Rao, S. Couet, G. S. Kar, K. Garello, and P. Gambardella, Tailoring the Switching Efficiency of Magnetic Tunnel Junctions by the Fieldlike Spin-Orbit Torque, Physical Review Applied 18, 044070 (2022b).
- Khvalkovskiy et al. [2013] A. V. Khvalkovskiy, D. Apalkov, S. Watts, R. Chepulskii, R. S. Beach, A. Ong, X. Tang, A. Driskill-Smith, W. H. Butler, P. B. Visscher, D. Lottis, E. Chen, V. Nikitin, and M. Krounbi, Basic principles of STT-MRAM cell operation in memory arrays, Journal of Physics D: Applied Physics 46, 074001 (2013).
- [27] See Supplemental Material at [URL will be inserted by publisher] for the magnetic characterization and the development of the LLG equation.
- Tillie et al. [2016] L. Tillie, E. Nowak, R. C. Sousa, M.-C. Cyrille, B. Delaet, T. Magis, A. Persico, J. Langer, B. Ocker, I.-L. Prejbeanu, and L. Perniola, Data retention extraction methodology for perpendicular STT-MRAM, in 2016 IEEE International Electron Devices Meeting (IEDM) (IEEE, San Francisco, CA, USA, 2016) pp. 27.3.1–27.3.4.
- Feng and Visscher [2004] X. Feng and P. B. Visscher, Sweep-rate-dependent coercivity simulation of FePt particle arrays, Journal of Applied Physics 95, 7043 (2004).
- Taniguchi et al. [2015] T. Taniguchi, S. Mitani, and M. Hayashi, Critical current destabilizing perpendicular magnetization by the spin Hall effect, Physical Review B 92, 024428 (2015).
- Torrejon et al. [2015] J. Torrejon, F. Garcia-Sanchez, T. Taniguchi, J. Sinha, S. Mitani, J.-V. Kim, and M. Hayashi, Current-driven asymmetric magnetization switching in perpendicularly magnetized CoFeB/MgO heterostructures, Physical Review B 91, 214434 (2015).
- Fan et al. [2019] W. Fan, J. Zhao, M. Tang, H. Chen, H. Yang, W. Lü, Z. Shi, and X. Qiu, Asymmetric Spin-Orbit-Torque-Induced Magnetization Switching With a Noncollinear In-Plane Assisting Magnetic Field, Physical Review Applied 11, 034018 (2019).
- Hao and Xiao [2015] Q. Hao and G. Xiao, Giant spin Hall effect and magnetotransport in a Ta/CoFeB/MgO layered structure: A temperature dependence study, Physical Review B 91, 224413 (2015).
- Mikuszeit et al. [2015] N. Mikuszeit, O. Boulle, I. M. Miron, K. Garello, P. Gambardella, G. Gaudin, and L. D. Buda-Prejbeanu, Spin-orbit torque driven chiral magnetization reversal in ultrathin nanostructures, Physical Review B 92, 144424 (2015).
- Baumgartner et al. [2017] M. Baumgartner, K. Garello, J. Mendil, C. O. Avci, E. Grimaldi, C. Murer, J. Feng, M. Gabureac, C. Stamm, Y. Acremann, S. Finizio, S. Wintz, J. Raabe, and P. Gambardella, Spatially and time-resolved magnetization dynamics driven by spin–orbit torques, Nature Nanotechnology 12, 980 (2017).
- Soucaille et al. [2016] R. Soucaille, M. Belmeguenai, J. Torrejon, J.-V. Kim, T. Devolder, Y. Roussigné, S.-M. Chérif, A. A. Stashkevich, M. Hayashi, and J.-P. Adam, Probing the Dzyaloshinskii-Moriya interaction in CoFeB ultrathin films using domain wall creep and Brillouin light spectroscopy, Physical Review B 94, 104431 (2016).
- Vigier [2023] J. Vigier, Writing determinism in SOT-MRAM memory bitcells with perpendicular anisotropy, Ph.D. thesis, Université Grenoble Alpes (2023).
Supplementary material

### v3

@@ IV Discussion and Simulations
To go further in the analysis of these results, we compared them with the results obtained from simulations. It is known that for pillar diameters such as those we have measured, the reversal of magnetization by SOT occurs through a two-step nucleation-propagation process [mikuszeit_spin-orbit_2015, baumgartner_spatially_2017]. Micromagnetic simulations are therefore the most accurate way of faithfully reproducing magnetization behavior. However, such simulations to calculate WER curves are time-consuming. Therefore, we aimed to reproduce these measurements using macrospin simulations. In this coherent reversal approximation, the magnetization dynamics can be reproduced by the Landau-Lifshitz-Gilbert equation, to which the two terms describing the FL and DL torques of the SOTs [park_macrospin_2014] are added:
@@
is obtained by Vibrating Sample Magnetometer measurements, neglecting the effect of any dead layer. , , and are imposed. is initially taken to be 0.033 [soucaille_probing_2016, vigier2023] and and H are first roughly estimated by fitting the SOT current vs H forward switching curve using the equation established in [lee_threshold_2013]:
@@
| Rise/Fall Time | 70 ps |
The comparison of the results of the numerical simulations and those of the experiments is shown in Fig. 3 (d), (e) and (f). The simulations do not reproduce all the characteristics of the experimental curves, for example, the variation of the switching current as a function of H field, which is very abrupt in the simulations compared with the behavior obtained experimentally. Nevertheless, the general agreement is very good, and the different regions of no-switching, deterministic switching with minimal WER, and loss of determinism in the switching, are reproduced very well. Furthermore, the parameter values used to obtain this agreement are realistic and very close to those reported in the literature for W/CoFeB/MgO multilayers (see for example [shao_roadmap_2021] and references herein). This is particularly noteworthy given that, as explained above, the dynamics of magnetization reversal by SOT can only be described in detail using a micromagnetic model.
This agreement between experimental and simulated results could point to the predominant role of nucleation in determining threshold currents for the boundaries between no-switching/deterministic switching and deterministic switching/back-switching for sub- nm SOT-MRAM devices [krizakova_tailoring_2022].
@@
In the presence of an applied in-plane field, H, the two equilibrium positions m are shifted towards the x-direction such that their new coordinates can be given by m and m where , as explained in [taniguchi_theoretical_2019]. In our simulations for I and H, the initial state of magnetization is m-. During the application of the current pulse, the magnetization reaches a state of dynamic equilibrium under the predominant action of SOT, the applied magnetic field and the anisotropy. At the end of the pulse, the magnetization relaxes under the effect of the applied magnetic field, the anisotropy and thermal fluctuations.
@@
When the current is increased, in the region of deterministic magnetization reversal, the current-induced torques are strong enough to drive the magnetization to a dynamic equilibrium close to m+ with small negative x and positive y components. At the end of the current pulse, the magnetization then relaxes to the m+ state, resulting in a deterministic switching, as shown in Fig. 4 (a) for IA and H Oe (green curve). By further increasing the current, the dynamic equilibrium begins to shift towards a fully in-plane position. This position is energetically unstable, and the resulting final state after the relaxation is highly sensitive to any energy or time fluctuation, as reported in [taniguchi_theoretical_2019]. As shown in Fig. 4 (a) (blue curve) for IA and H Oe, the magnetization can then return to its original m- state, resulting in a BSW phenomenon.
@@
With the same arguments, we can infer the impact of H, and consequently H, by comparing the switching orbits for IA and H Oe in the presence of a magnetic field H Oe and Oe as shown in Fig. 4 (c). By expanding equation 1, we obtained the torque components due to the current-induced SOT () and externally applied HY (). The z-component of is given by equation 3.
With the parameter values given in Table 1, the second term in equation 3 is dominated by the current-induced term, , with negligible effect of . Whereas in the first term in equation 3, and are comparable and compete. Under a positive injected current both and mX are negative resulting in overall positive out-of-plane torques pulling the magnetization towards m. Starting from H (Fig. 4 (a) blue curves), the application of H (H Oe) opposes the z-component of the current-induced torques, reducing the out-of-plane torques. This reduction in the z-component of the torques further shifts the dynamic equilibrium in-plane, reinforcing the non-deterministic character of the reversal, as shown by the yellow curve in Fig. 4 (c). On the other hand, increasing the z-component of the torques pulls the dynamic equilibrium away from this energetically unstable position. Consequently, the application of H Oe leads to a deterministic reversal, as shown by the brown curve in Fig. 4 (c). Further details are provided in [supplementary]. This discussion also provides insights into the impact of H () since it is modulated by H. It should be noted that these results are in line with previous studies showing deterministic switching for and non-deterministic switching for [lee_oscillatory_2018, yoon_anomalous_2017, krizakova_tailoring_2022]. In the second case, the two terms in equation 3 are opposite in sign and compete to further reduce the out-of-plane torque component, contributing to shifting the dynamic equilibrium towards an in-plane position. Although previous studies [lee_oscillatory_2018, yoon_anomalous_2017, krizakova_tailoring_2022] considered the competition between H and H for switching determinism, our discussion above highlights the competition between H (or ) and H. Since in Ta-based devices [lee_oscillatory_2018, yoon_anomalous_2017, krizakova_tailoring_2022], neglecting the term is reasonable and results in identical discussions.
@@
Since the final state depends strongly on the relaxation process, it should also depend on the magnetic damping parameter as reported in [lee_threshold_2013, park_macrospin_2014, taniguchi_reduction_2020]. Fig. 4 (d) shows the switching orbits for IA, H Oe and different . The damping coefficient , which we used in our simulations, results in slow energy dissipation and hence randomness in the final state (blue curve). This randomness is even more prevalent when is lower, as shown by the light blue curve for . As increases, energy is dissipated more rapidly and randomness is reduced, eventually disappearing, as shown by the red curve for and as reported previously [taniguchi_reduction_2020]. Unlike STT switching current, which scales with , SOT switching current is little affected by a large magnetic damping parameter [lee_threshold_2013, taniguchi_critical_2015]. Increasing the magnetic damping parameter can therefore significantly reduce the BSW without severely affecting the SOT-MRAM write current. Finally, since thermal fluctuations introduce randomness in the current-induced torques during the relaxation process, higher temperatures show an increase in WER. At low temperature, the thermal fluctuations are minimal and the final state depends only on the dynamic equilibrium and .
@@ V Reducing the backswitching by shaping the current pulse
There is another way of reducing BSW and improving WER without having to change the material system. It consists of increasing the length of the fall time of the write current pulse, as predicted numerically [zhu_threshold_2020]. Fig. 5 (a) shows WER curves obtained experimentally in the presence of an in-plane field, H Oe and injecting 10 ns-long SOT current pulses. A widening of the forward switching region is observed as the fall time increases from 2 ns to 4 ns, with the BSW phenomenon disappearing for the current range used in the case of a fall time of 4 ns. This can be understood using the same arguments as above, by looking at magnetization relaxation at the end of the pulse. When the fall time is non-zero, current-induced torques are present during relaxation. The longer the fall time, the longer their amplitude remains at high values. The impact of thermal fluctuations is then reduced, but more importantly, the magnetization is gradually brought out of an energetically unstable position. In this way, the dynamic equilibrium is gradually modified. Such behavior is in line with the explanation given in [zhu_threshold_2020]. The effect of this increase in fall time duration in the presence of a transverse field H or an out-of-plane field H has not been experimentally tested, but a discussion based on macrospin simulations can be found in [supplementary]. However, much longer fall times (4 ns) were required in our experiments compared with those obtained in the simulations of [zhu_threshold_2020], typically of the order of 0.2 ns. It may be attributed to thermal fluctuations and Joule heat effects, which are not taken into account in the simulations. This difference needs to be investigated. In any case, the results of our experiments show a simple solution to the BSW phenomenon that can be easily implemented without the need for further material developments. We tested this solution on a complete SOT-MRAM single cell (Fig. 5 (b)). We improved the WER by almost 3 orders of magnitude using a longer fall time of 2 ns, and obtained no error for write attempts.
@@ Acknowledgements.
We thank the staff of Antaios for their support throughout this project. We thank Chando Park, Minrui Yu, Luc Thomas and Mahendra Pakala from Applied Materials for fruitful discussions. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 955671, and was supported by the Région Auvergne-Rhône-Alpes Pack Ambition Recherche Program (grant 19-009938-01-MAPS) and the French RENATECH network implemented at the Upstream Technological Platform in Grenoble PTA (ANR-22-PEEL-0015).Supplementary material